OpenStudy (ttop0816):

am i correct?? The slope of a line is -1/3, and the y-intercept is 10/3. What is the equation of the line written in general form? 10x + 3y - 1 = 0 x + 3y + 10 = 0 x + 3y - 10 = 0 <<<

OpenStudy (kc_kennylau):

yes my friend you're correct :D

OpenStudy (luigi0210):

^ Like the wise owl said, yes

OpenStudy (kc_kennylau):

Since the slope if -1/3 and the y-intercept is 10/3, we can write the equation in slope-intercept form (y=mx+b where m is slope and b is y-intercept): \[y=-\frac13x+\frac{10}3\] Then, we can put everything to the left to from the first step of the general solution: \[y+\frac13x-\frac{10}3=0\] However, we can't have fractional co-efficient in the general form, so we have to multiply the whole equation by 3: (Math rules suck, right? :P) \[3y+x-10=0\] Then, we need to rearrange the terms (ditto): \[x+3y-10=0\]

OpenStudy (luigi0210):

Way to type bro

OpenStudy (ttop0816):

and also for this question am i correct?? Write the equation of the line that has an x-intercept of -3 and passes through the point (-3, 7). x = -3 y = -3 y = 7 <--